### v3

@@ 5.3 Players’ Performance
To ensure the accuracy of our measurement of players’ genuine attitudes toward the LGB community, it is crucial to assess whether shifting away from Graves to other characters impacts players’ performance. If there are performance costs, our estimates could be biased toward zero, as players might continue using Graves for strategic considerations. Moreover, if players switch characters primarily for convenience, our analysis might unintentionally capture a different phenomenon instead of the intended social stigma.
@@
We employ difference-in-differences identification and estimation strategies to assess the impact of players abandoning Graves on their performance. We gauge players’ performance by their daily win rate. Our analysis focuses on the $953$ prior-users of Section 5.2, who are classified into treated or control groups based on their responses to Graves’ disclosure. We consider two different definitions of the treatment, sorted by their intensity. In the first version, labeled moderate reduction, we classify as treated those players who decreased their average pick rate for Graves following his disclosure by at most $75\%$ of their pre-treatment average pick rate (the number of treated units is $287$). In the second version, labeled substantial reduction, we classify as treated those players who reduced their average pick rate for Graves by $75\%$ to $100\%$ post-disclosure (the number of treated units is 491491).343434 Figure A.VII in Appendix A shows the distribution of percentage reduction in average pick rates for Graves from pre-treatment to post-treatment among prior users. In both scenarios, the control group remains the same and consists of the $175$ prior users who did not reduce their average pick rate for Graves following his disclosure.
@@
Under the standard assumptions of parallel trends and no anticipation [[, see, e.g.,]]roth2023s, we can identify the average treatment effect on the treated (ATT) using observable data. The parallel trend assumption posits that the performance of treated and untreated players would have evolved similarly if Graves’ disclosure had not occurred. While we cannot formally test this assumption, the findings of Section 5.1 and Section 5.2 provide substantial support for its plausibility.353535 Moreover, we demonstrate below the absence of pre-treatment differences in trends by reporting placebo estimates of the ATT that are not statistically different from zero. This is often viewed as a natural plausibility check, although even if pre-trends are perfectly parallel, this does not necessarily guarantee the satisfaction of the post-treatment parallel trends assumption [[, see, e.g.,]]roth2023s. As for the no anticipation assumption, it stipulates that in the weeks preceding the disclosure, players’ performance did not change due to the incoming Graves’ disclosure. The plausibility of this assumption was thoroughly discussed in Section 2.2 and Section 4.1.
@@
We implement the approach of [181] to target the ATT at a particular day $t>T^{pre}$:363636 The framework outlined in [181] is broader as it accommodates multiple groups defined by the timing of treatment reception. This enables the identification and estimation of the group-time ATTs, defined as $ATT\left(g,t\right):=\operatorname{\mathbb{E}}\left[Y_{i,t}\left(g\right)-Y_{i,t}\left(0\right)|G_{g}=1\right]$, where $G_{g}$ is a binary variable indicating treatment reception in period $g$. However, our data set features a single group, given that all treated players receive the treatment simultaneously (i.e., at Graves’ disclosure date). This allows us to simplify notation and focus on the time ATTs in equation (6) for the single group we observe.
$ATT(t):=\operatorname{\mathbb{E}}[Y_{i,t}\left(1\right)-Y_{i,t}\left(0\right)|D_{i}=1],$ (6)
@@
where potential outcomes are defined as in Section 3.2, and $D_{i}$ is a binary variable indicating whether a player is treated or not. Under the assumptions of parallel trends and no anticipation, [181] show that $ATT(t)$ can be identified by comparing the change in outcomes between the latest period before the coming-out event and day $t$ experienced by treated players to the change in outcomes experienced by control players.373737 Formally, [181] show that $ATT(t)=\operatorname{\mathbb{E}}[Y_{i,t}-Y_{i,T^{pre}}|D_{i}=1]-\operatorname{\mathbb{E}}[Y_{i,t}-Y_{i,T^{pre}}|D_{i}=0]$. Estimation is carried out by replacing expectations with their sample analogs.
@@
Figure 7: Point estimates and simultaneous $95\%$ confidence bands allowing for clustering at the player level for the $ATT(t)$. Red lines refer to pre-treatment periods, while blue lines refer to post-treatment periods. Each row corresponds to a different version of the treatment.
@@
Figure 7 displays the point estimates and simultaneous $95\%$ confidence bands for the $ATT(t)$. Overall, we find that shifting away from Graves to other characters has no impact on players’ performance. None of the estimated $ATT(t)$ is statistically different from zero, suggesting that transitioning to other characters does not result in any performance-related consequences. This finding highlights that the decision to move away from Graves is not influenced by performance considerations.
@@
Figure 7 further displays placebo estimates of the time ATTs for the ten days before the treatment.383838 Figure A.VIII in Appendix A displays the remaining estimated placebo $ATT(t)$. As explained above, these estimates are valuable for “pre-testing” the credibility of the parallel trend assumption [181]. Notably, all placebo time ATTs in the pre-treatment periods are statistically insignificant, supporting the validity of the parallel trends assumption.
5.4 New Substitute Character
@@
On June $9^{th}$, $2022$, a new character (Bel’Veth) was released.393939 Figure A.IX in Appendix A displays Bel’Veth’s daily pick rates. Since the primary position the character is designed for is the same position as Graves is designed for, it can be considered a close substitute. Therefore, players’ decisions to switch away from Graves (after June $9^{th}$) might to some degree be driven by the desire to experiment with the new character and to explore potential competitive advantages, challenging social stigma as the primary explanation behind our main result.
@@
Figure 8: Average pick rate for Bel’Veth after its release among prior users.
@@
If the release serves as the primary explanation for the observed drop in Graves pick rate we would expect a positive correlation between the size of the drop in players’ Graves usage and players preferences for the new character. In Figure 8 we show that this is not the case. Figure 8 shows the average pick rate for Bel’Veth for the prior users of Section 5.2, classified as in Section 5.3. We find that those players who feature a substantial reduction in their Graves’ pick rates following the coming-out event are less likely to pick Bel’Veth for their matches than players who feature a moderate reduction. Moreover, these players are even less likely to pick Bel’Veth for their matches than players who did not react at all to Graves’ disclosure. We therefore argue that the release is unlikely to serve as the primary explanation for the observed main result of the paper.
5.5 Coming Out versus LGBT Pride Month
@@
In Appendix D, we introduce a theoretical framework that formalizes the existence of two simultaneous treatments and discuss the implications for interpretation. Additionally, we outline sufficient assumptions that enable us to separate the impacts of coming out and LGBT Pride Month on players’ preferences for Graves. Here, we provide the main intuitions behind our approach, directing the reader to the appendix for technical details.
@@
Figure 9: Composite LGB unit’s daily pick rates and synthetic control estimation results. The actual series is smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
@@
We create a composite LGB unit by averaging the pick rates of Diana, Leona, Nami, and Neeko and employ the synthetic control estimator described in Section 3.2 to estimate the effect of LGBT Pride Month on players’ preferences for LGB characters. Then, under the assumption that the influence of LGBT Pride Month is uniform across all LGB characters, we can compare the results with those obtained for Graves to separate the impacts of coming out and LGBT Pride Month on players’ preferences for Graves. Intuitively, if the estimated impact of LGBT Pride Month on players’ preferences for LGB characters is small relative to the estimated impact of the coming-out event on players’ preferences for Graves, this suggests that the findings of Section 4 must be primarily attributed to Graves’ disclosure.
@@
Figure 9 displays the actual and the synthetic pick rate series for the composite LGB unit. Overall, our analysis suggests that LGBT Pride Month had no impact on players’ preferences for LGB characters. Before the treatment, the synthetic control estimator closely aligns with the actual series, providing support for the estimator’s ability to predict the counterfactual series. After the treatment date, the synthetic control estimator continues to align with the actual series, confirming that the players’ preferences for LGB characters were unaffected by LGBT Pride Month. The average effect is estimated to be 0.215-0.215 percentage points (standard error: $3.052$), and the conventional $95\%$ confidence interval encompasses zero, indicating a failure to reject the null hypothesis of no effect. Under the homogeneity assumption discussed above, these findings support the interpretation that the estimated effects presented in Section 4 are primarily driven by Graves’ disclosure rather than being influenced by the broader context of LGBT Pride Month.
6 Conclusion
Discrimination based on sexual orientation is first and foremost a human rights issue. However, when individuals with a stigmatized identity are unfairly targeted in education, health, social, and political settings, there is a loss of human capital that can have detrimental effects on the economy as a whole [174]. For example, bullying and discrimination act as barriers to students’ acquisition of skills and knowledge. Furthermore, even short experiences of bullying can have severe long-term health consequences [160]. Therefore, understanding the barriers that individuals from stigmatized groups face is of large societal importance.
In this study, we utilize a comprehensive data set sourced from the widely popular online video game League of Legends and exploit exogenous variation in the identity of a playable character to credibly identify sentiments towards LGB status. Players in the game select a playable character before each match. Each playable character is characterized by game-relevant attributes and a background story. Leveraging an unexpected revelation during the 2022 LGBT Pride Month, wherein game developers disclosed the sexual orientation minority status of a playable character, we investigate individuals’ responses to this disclosure. By tracking players’ revealed preferences for the character over a meaningful period, we provide insights into reactions following the disclosure. To isolate the effects on player preferences from potential confounding influences, we employ synthetic control methods. Our findings reveal a substantial and persistent negative impact, with preferences for this character decreasing by more than 30% over a meaningful period. This underscores the potential negative consequences of disclosing one’s sexual minority status and provides a rationale for the underrepresentation of individuals with LGB status in many regions and professions.
To bolster the credibility of stigma as the primary explanation for the estimated effects, we address and eliminate several alternative channels. First, we rule out the possibility that shifts in characters’ relative strengths could explain our estimated effect. Second, we show that players’ skills have no correlation with the choice to drop the character. Third, we provide evidence that switching to other characters does not affect the performance of the players involved and that the release of new characters in the post-treatment period is unlikely to serve as primary explanation behind the results. Fourth, we introduce a theoretical framework that formalizes the existence of two ”simultaneous treatments” and use information on other playable characters that belong to the LGB group to rule out that LGBT Pride Month serves as the main explanation for the observed result.
We have dismissed the possibility that any actual or presumed change in the character’s strength is driving our results. This strongly suggests that the estimated cost of coming out is unlikely to be driven by factors other than stigmatization. This insight holds significant implications for policymakers aiming to develop interventions that effectively tackle stigmatization and improve the overall well-being of LGB individuals. Policies should be formulated to discourage discriminatory behavior, either by increasing its costs or by creating inclusive social environments that promote the acceptance of sexual minority individuals and reduce the stigma. Raising awareness about the reaction to sexual minority disclosure could be an important step to develop such a society, as previous research has demonstrated the value of discrimination awareness for minority groups outcomes [170]. At the same time, policymakers may also consider providing resources and support to individuals who have recently come out, such as access to counseling and mental health services. By doing so, they can mitigate some of the negative outcomes that may arise from coming out.
Appendix A Additional Figures and Tables
Figure A.I: Histograms of players’ total and average daily activity, measured in both number of matches played and hours spent in-game.
Figure A.II: Identities and contributions of characters in the donor pool for the Graves’ synthetic control displayed in Figure 2.
Figure A.III: Ratio of post- to pre-treatment RMSE for Graves and control characters. Only control characters with a pre-treatment RMSE of at least 1 are considered to exclude overfit specifications.
Figure A.IV: Distribution of overall win rates and total matches played for players classified as above or below the median pre-treatment win rate.
Figure A.V: Identities and contributions of characters in the donor pool for the Graves’ synthetic controls displayed in Figure 4.
Figure A.VI: Shares of matches played in each role before and after the coming-out event. Players are divided into two groups based on their preferences for Graves before his disclosure.
Figure A.VII: Distribution of percentage reduction in average pick rates for Graves from pre-treatment to post-treatment among prior users.
Figure A.VIII: Point estimates and simultaneous $95\%$ confidence bands allowing for clustering at the player level for the placebo $ATT(t)$. Each row corresponds to a different version of the treatment.
Figure A.IX: Bel’Veth’s daily pick rates. The series is smoothed by a Nadaraya-Watson regression. The rainbow area highlights LGBT Pride Month. The non-zero pick rate displayed in early January is a result of the smoothing.
Table A.I: Most popular characters.
Table A.II: Regional results.
Synthetic Controls Regularized Synthetic Controls (1) (2) (3) (4) All characters Only non-substitutes All characters Only non-substitutes Panel 1: Europe $\hat{\tau} -$9.062 -11.719 -8.696 -10.670 $95\%$ CI [-15.832, -2.292] [-19.479, -3.958] [-14.210, -3.183] [-19.594, -1.758] N. Donors 5 5 6 8 RMSE 2.206 2.639 2.155 2.583 Pre-treatment average 14.618 14.618 14.618 14.618 Panel 2: Korea $\hat{\tau} -$10.015 -8.955 -10.299 -8.990 $95\%$ CI [-18.262, -1.768] [-20.502, 2.593] [-17.720, -2.878] [-18.368, 0.376] N. Donors 3 3 3 3 RMSE 6.211 6.550 6.249 6.594 Pre-treatment average 31.441 31.441 31.441 31.441 Panel 3: Latin America $\hat{\tau} -$6.212 -5.899 -5.206 -4.960 $95\%$ CI [-11.850, -0.574] [-12.449, 0.651] [-10.645, 0.233] [-12.531, 2.605] N. Donors 4 4 7 6 RMSE 2.575 2.552 2.514 2.522 Pre-treatment average 16.217 16.217 16.217 16.217 Panel 4: North America $\hat{\tau}$ 0.939 0.940 1.110 1.130 $95\%$ CI [ -4.583, 6.462] [ -6.259, 8.138] [ -3.412, 5.632] [ -5.807, 8.079] N. Donors 3 3 4 4 RMSE 4.669 4.669 4.681 4.685 Pre-treatment average 22.282 22.282 22.282 22.282
Notes. Point estimates and $95\%$ confidence intervals for $\hat{\tau}$. Additionally, the number of donors receiving a non-zero weight and the pre-treatment root mean squared error are displayed. Each panel reports the results obtained using only matches from a particular macro-region. Each column corresponds to a different specification, with the specifications differing solely in the employed estimator and donor pool composition.
Appendix B Robustness Checks
This section presents a series of robustness checks that support the credibility of our findings. We begin with checks related to our main results, followed by additional analyses that validate the player heterogeneity findings.
B.1 Main Results
We examine the robustness of our main results from Section 4.1 to alternative estimation strategies and variations in the composition of the donor pool. In particular, we repeat our analysis employing a regularized synthetic control estimator (see Section 3.2) and explore different donor pool configurations focusing on characters from distinct roles. Notably, Graves is predominantly designed for and played in three of the possible roles within a team. Consequently, there is a possibility of spillover effects on other characters mainly played in these positions, as players transitioning away from Graves are likely to switch to these alternatives.404040 Table 1 and the findings of Section 5.2 support this intuition. To mitigate this potential for spillover effects, we restrict our donor pool to characters that are “non-substitutes” of Graves, that is, those primarily designed for the remaining two roles.414141 Graves is predominantly designed for and played in the top, jungle, and mid positions. Therefore, we consider characters designed for and played in the bottom and support positions as “non-substitutes.”
Table B.I: Main results.
Synthetic Controls Regularized Synthetic Controls (1) (2) (3) (4) All characters Only non-substitutes All characters Only non-substitutes $\hat{\tau} -$7.606 -7.161 -7.323 -7.080 $95\%$ CI [-13.542, -1.670] [-15.383, 1.062] [-11.471, -3.174] [-14.831, 0.665] N. Donors 4 4 4 5 RMSE 2.365 2.525 2.330 2.388 Pre-treatment average 18.558 18.558 18.558 18.558
Notes. Point estimates and $95\%$ confidence intervals for $\hat{\tau}$. Additionally, the number of donors receiving a non-zero weight and the pre-treatment root mean squared error are displayed. Each column corresponds to a different specification, with the specifications differing solely in the employed estimator and donor pool composition.
Table B.I displays the results. For any donor pool composition, the results are not sensitive to the choice of the regularization parameter, and point estimates are consistently negative. The results are consistent also quantitatively across all specifications, with a decline ranging between $40.98\%$ and $38.17\%$ of the pre-treatment average preferences for Graves. Overall, these results support our main finding of a substantial negative impact of the coming-out event on players’ preferences for Graves.
Figure B.I: Robustness checks for main results. The upper panel shifts the coming-out event ten days earlier, with the new treatment date denoted by the vertical gray dashed line. The lower panel reports leave-one-out estimates of the synthetic control series, obtained by excluding one of the characters of Figure A.II at a time from the donor pool.
We also assess the credibility of the synthetic control estimator by conducting a robustness check that artificially shifts the coming-out event ten days earlier. This backdating exercise allows us to evaluate the estimator’s predictive accuracy during a ten-day hold-out period [[, see e.g.,]]abadie2022synthetic. The upper panel of Figure B.I presents the results of this analysis. We observe three key findings. First, the estimated effects remain qualitatively and quantitatively consistent, confirming a negative and persistent impact of the coming-out event on players’ revealed preferences for Graves. Second, the synthetic control estimator demonstrates a good fit during the hold-out period, indicating its ability to accurately capture Graves’ behavior prior to the disclosure. Third, the actual and the synthetic series begin to diverge on the true day of disclosure, even when the estimator has no knowledge of the actual disclosure date. The absence of estimated effects before the coming-out event also lends support to the plausibility of a no-anticipation assumption [[, see e.g.,]]abadie2021using.
Finally, we conduct an additional robustness test by performing a leave-one-out exercise, where we repeatedly estimate the synthetic control series by excluding one character with non-zero estimated weights at a time from the donor pool [[, see e.g.,]]abadie2021using. The lower panel of Figure B.I presents the results of this analysis. Overall, our finding of a negative and persistent impact of the coming-out event on players’ preferences for Graves is robust to the exclusion of any particular character. Most of the leave-one-out synthetic series closely align with the main estimate, thus reinforcing the robustness of the main conclusion of our study. One leave-one-out series falls beneath the other synthetic series, suggesting a somewhat reduced, although still negative, impact. However, this series diverges from the actual series in the weeks prior to the treatment, which undermines the reliability of its results.424242 This series is obtained by excluding the character Ezreal from the donor pool. Figure A.II in Appendix A shows that this character receives the largest weight in our main specification. Therefore, the divergence of this leave-one-out series from the actual series is unsurprising.
B.2 Player Heterogeneity
We assess the robustness of our player heterogeneity results from Section 4.2 using the same backdating and leave-one-out exercises described in the previous subsection. Figure B.II displays the results. For both skill groups, we observe that even when the estimator has no knowledge of the actual disclosure date, the actual and synthetic series begin to diverge precisely on the true day of disclosure. The fit during the hold-out period is also strong, indicating good predictive performance. Moreover, most of the leave-one-out synthetic series closely track the main estimates, reinforcing the robustness of our findings.
(a) Players below median skill
(b) Players above median skill
Figure B.II: Robustness checks for player heterogeneity results. The upper panels shift the coming-out event ten days earlier, with the new treatment date denoted by the vertical gray dashed line. The lower panels report leave-one-out estimates of the synthetic control series, obtained by excluding one character with non-zero estimated weights at a time from the donor pool.
Appendix C Twisted Fate
@@
In the League of Legends universe, Graves, described as a “gruff-looking, broad-shouldered, middle-aged man,” forms a criminal partnership with Twisted Fate, who is characterized as “a tall, handsome male with tanned skin, trimmed beard, and long dark hair.” Together, they engage in various illicit activities until, in the midst of a heist, Graves finds himself captured. This event leads Graves to perceive a betrayal by Twisted Fate, prompting a pursuit of revenge upon his escape. However, the duo ultimately decides to reconcile their differences and resumes their collaboration.434343 The complete background of Graves and Twisted Fate is available at https://leagueoflegends.fandom.com/wiki/Graves and https://leagueoflegends.fandom.com/wiki/Twisted_Fate.
@@
We investigate whether this implied revelation has captured the players’ attention in Figure C.I, illustrating the Google search interest for the queries “Graves gay” and “Twisted Fate gay.” Throughout $2022$, we observe approximately no interest in the latter query, with a small spike occurring during the week of the coming-out event, amounting to less than half of the spike associated with Graves. Furthermore, the search interest for Twisted Fate is always lower than that for Graves, suggesting the greater popularity of Graves among players.444444 This is also suggested by Graves’s pick rates being approximately 3 to 4 times higher than those of Twisted Fate, as displayed in Figure C.II. These results underscore the relatively low attention directed towards Twisted Fate from players, who were primarily focused on Graves and the explicit establishment of his sexual orientation. As a result, we concentrate our analysis on Graves and his disclosure for a more credible identification of the effects of coming out.
Figure C.I: Google search interest over time for the queries “Graves gay” and “Twisted Fate gay.” The dashed vertical line denotes the week of disclosure, and the shaded area highlights the League of Legends World Championship.
Figure C.II: Graves and Twisted Fate’s daily pick rates. The series are smoothed by a Nadaraya-Watson regression. The rainbow area highlights LGBT Pride Month.
Appendix D Anatomy of the Coming-Out Event
In this section, we discuss how the existence of two treatments - the disclosure of Graves’ sexual orientation and the start of LGBT Pride Month - occurring at the same time may affect the interpretation of the main findings of Section 4.1. The notation follows that used in Section 3.2. The results of the analysis are detailed in Section 5.5.

### v4

@@ D.3 Players’ Performance
To ensure that our estimates capture behavioral responses to the disclosure rather than strategic considerations, it is crucial to assess whether shifting away from Graves to other characters impacts players’ performance. If there are performance costs, our estimates could be biased toward zero, as players might continue using Graves for strategic considerations. Moreover, if players switch characters primarily for convenience or strategic reasons, our analysis might capture a different phenomenon unrelated to the disclosure.
@@
We employ difference-in-differences identification and estimation strategies to assess the impact of players abandoning Graves on their performance. We gauge players’ performance by their daily win rate. Our analysis focuses on the $953$ prior-users of Section 5.2, who are classified into treated or control groups based on their responses to Graves’ disclosure. We consider two different definitions of the treatment, sorted by their intensity. In the first version, labeled moderate reduction, we classify as treated those players who decreased their average pick rate for Graves following his disclosure by at most $10$ percentage points (the number of treated units is $522$). In the second version, labeled substantial reduction, we classify as treated those players who reduced their average pick rate for Graves by $10$ to $100$ percentage points (the number of treated units is 256256).434343 Figure A.VIII in Supplementary Information A shows the distribution of changes in average Graves’ pick rates between the pre- and post-disclosure periods among prior users. In both scenarios, the control group remains the same and consists of the $175$ prior users who did not reduce their average pick rate for Graves following his disclosure.
@@
Under the standard assumptions of parallel trends and no anticipation [[, see, e.g.,]]roth2023s, we can identify the average treatment effect on the treated (ATT) using observable data. The parallel trend assumption posits that the performance of treated and untreated players would have evolved similarly if Graves’ disclosure had not occurred. While we cannot formally test this assumption, the findings of Section 5.1 and Section 5.2 provide substantial support for its plausibility.444444 Moreover, we demonstrate below the absence of pre-treatment differences in trends by reporting placebo estimates of the ATT that are not statistically different from zero. This is often viewed as a natural plausibility check, although even if pre-trends are perfectly parallel, this does not necessarily guarantee the satisfaction of the post-treatment parallel trends assumption [[, see, e.g.,]]roth2023s. As for the no anticipation assumption, it stipulates that in the weeks preceding the disclosure, players’ performance did not change due to the incoming Graves’ disclosure. The plausibility of this assumption was thoroughly discussed in Section 2.2 and Section 4.1.
@@
We implement the approach of [19] to target the ATT at a particular day $t>T^{pre}$:454545 The framework outlined in [19] is broader as it accommodates multiple groups defined by the timing of treatment reception. This enables the identification and estimation of the group-time ATTs, defined as $ATT\left(g,t\right):=\operatorname{\mathbb{E}}\left[Y_{i,t}\left(g\right)-Y_{i,t}\left(0\right)|G_{g}=1\right]$, where $G_{g}$ is a binary variable indicating treatment reception in period $g$. However, our data set features a single group, given that all treated players receive the treatment simultaneously (i.e., at Graves’ disclosure date). This allows us to simplify notation and focus on the time ATTs in equation (D.1) for the single group we observe.
$ATT(t):=\operatorname{\mathbb{E}}[Y_{i,t}\left(1\right)-Y_{i,t}\left(0\right)|D_{i}=1],$ (D.1)
@@
where potential outcomes are defined as in Section B, and $D_{i}$ is a binary variable indicating whether a player is treated or not. Under the assumptions of parallel trends and no anticipation, [19] show that $ATT(t)$ can be identified by comparing the change in outcomes between the latest period before the coming-out event and day $t$ experienced by treated players to the change in outcomes experienced by control players.464646 Formally, [19] show that $ATT(t)=\operatorname{\mathbb{E}}[Y_{i,t}-Y_{i,T^{pre}}|D_{i}=1]-\operatorname{\mathbb{E}}[Y_{i,t}-Y_{i,T^{pre}}|D_{i}=0]$. Estimation is carried out by replacing expectations with their sample analogs.
@@
Figure D.III: Point estimates and simultaneous $95\%$ confidence bands allowing for clustering at the player level for the $ATT(t)$. Red lines refer to pre-treatment periods, while blue lines refer to post-treatment periods. Each row corresponds to a different version of the treatment.
@@
Figure D.III displays the point estimates and simultaneous $95\%$ confidence bands for the $ATT(t)$. Overall, we find that shifting away from Graves to other characters has no impact on players’ performance. None of the estimated $ATT(t)$ is statistically different from zero, suggesting that transitioning to other characters does not result in any performance-related consequences. This finding highlights that the decision to move away from Graves is not influenced by performance considerations.
@@
Figure D.III further displays placebo estimates of the time ATTs for the ten days before the treatment.474747 Figure A.IX in Supplementary Information A displays the remaining estimated placebo $ATT(t)$. As explained above, these estimates are valuable for “pre-testing” the credibility of the parallel trend assumption [19]. Notably, all placebo time ATTs in the pre-treatment periods are statistically insignificant, supporting the validity of the parallel trends assumption.
D.4 New Substitute Character
@@
On June $9^{th}$, $2022$, a new character (Bel’Veth) was released.484848 Figure A.X in Supplementary Information A displays Bel’Veth’s daily pick rates. Since the primary position the character is designed for is the same position as Graves is designed for, it can be considered a close substitute. Therefore, players’ decisions to switch away from Graves (after June $9^{th}$) might to some degree be driven by the desire to experiment with the new character and to explore potential competitive advantages, challenging social stigma as the primary explanation behind our main result.
@@
Figure D.IV: Average pick rate for Bel’Veth after its release among prior users.
@@
If the release serves as the primary explanation for the observed drop in Graves pick rate we would expect a positive correlation between the size of the drop in players’ Graves usage and players preferences for the new character. In Figure D.IV we show that this is not the case. Figure D.IV shows the average pick rate for Bel’Veth for the prior users of Section 5.2, classified as in Section 5.3. We find that those players who feature a substantial reduction in their Graves’ pick rates following the coming-out event are less likely to pick Bel’Veth for their matches than players who feature a moderate reduction. Moreover, these players are even less likely to pick Bel’Veth for their matches than players who did not react at all to Graves’ disclosure. We therefore argue that the release is unlikely to serve as the primary explanation for the observed main result of the paper.
D.5 Coming Out versus LGBT Pride Month
@@
In Supplementary Information F, we introduce a theoretical framework that formalizes the existence of two simultaneous treatments and discuss the implications for interpretation. Additionally, we outline sufficient assumptions that enable us to separate the impacts of coming out and LGBT Pride Month on players’ preferences for Graves. Here, we provide the main intuitions behind our approach, directing the reader to the supplementary information for technical details.
@@
Figure D.V: Composite LGB unit’s daily pick rates and synthetic control estimation results. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
@@
We create a composite LGB unit by averaging the pick rates of Diana, Leona, Nami, and Neeko and employ the synthetic control estimator described in Section B to estimate the effect of LGBT Pride Month on players’ preferences for LGB characters. Then, under the assumption that the influence of LGBT Pride Month is uniform across all LGB characters, we can compare the results with those obtained for Graves to separate the impacts of coming out and LGBT Pride Month on players’ preferences for Graves. Intuitively, if the estimated impact of LGBT Pride Month on players’ preferences for LGB characters is small relative to the estimated impact of the coming-out event on players’ preferences for Graves, this suggests that the findings of Section 4 must be primarily attributed to Graves’ disclosure.
@@
Figure D.V displays the actual and the synthetic pick rate series for the composite LGB unit. Overall, our analysis suggests that LGBT Pride Month had no impact on players’ preferences for LGB characters. Before the treatment, the synthetic control estimator closely aligns with the actual series, providing support for the estimator’s ability to predict the counterfactual series. After the treatment date, the synthetic control estimator continues to align with the actual series, confirming that the players’ preferences for LGB characters were unaffected by LGBT Pride Month. The average effect is estimated to be 0.672-0.672 percentage points (standard error: $2.806$), and the conventional $95\%$ confidence interval encompasses zero, indicating a failure to reject the null hypothesis of no effect. Under the homogeneity assumption discussed above, these findings support the interpretation that the estimated effects presented in Section 4 are primarily driven by Graves’ disclosure rather than being influenced by the broader context of LGBT Pride Month.
Supplementary Information E Twisted Fate
@@
In the League of Legends universe, Graves, described as a “gruff-looking, broad-shouldered, middle-aged man,” forms a criminal partnership with Twisted Fate, who is characterized as “a tall, handsome male with tanned skin, trimmed beard, and long dark hair.” Together, they engage in various illicit activities until, in the midst of a heist, Graves finds himself captured. This event leads Graves to perceive a betrayal by Twisted Fate, prompting a pursuit of revenge upon his escape. However, the duo ultimately decides to reconcile their differences and resumes their collaboration.494949 The complete background of Graves and Twisted Fate is available at https://leagueoflegends.fandom.com/wiki/Graves and https://leagueoflegends.fandom.com/wiki/Twisted_Fate.
@@
We investigate whether this implied revelation has captured the players’ attention in Figure E.I, illustrating the Google search interest for the queries “Graves gay” and “Twisted Fate gay.” Throughout $2022$, we observe approximately no interest in the latter query, with a small spike occurring during the week of the coming-out event, amounting to less than half of the spike associated with Graves. Furthermore, the search interest for Twisted Fate is always lower than that for Graves, suggesting the greater popularity of Graves among players.505050 This is also suggested by Graves’s pick rates being approximately 3 to 4 times higher than those of Twisted Fate, as displayed in Figure E.II. These results underscore the relatively low attention directed towards Twisted Fate from players, who were primarily focused on Graves and the explicit establishment of his sexual orientation. As a result, we concentrate our analysis on Graves and his disclosure for a more credible identification of the effects of coming out.
Figure E.I: Google search interest over time for the queries “Graves gay” and “Twisted Fate gay.” The dashed vertical line denotes the week of disclosure, and the shaded area highlights the League of Legends World Championship.
Figure E.II: Graves and Twisted Fate’s daily pick rates. The rainbow area highlights LGBT Pride Month.
Supplementary Information F Anatomy of the Coming-Out Event
In this section, we discuss how the existence of two treatments—the disclosure of Graves’ sexual orientation and the start of LGBT Pride Month—occurring at the same time may affect the interpretation of the main findings of Section 4.1. The notation follows that used in Section B. The results of the analysis are detailed in Section 5.5.